Maria needs to mix a 10% alcohol solution with a 50% solution to create a 100 millimeters of a 34% solution how many millileters of each solution must Maria use
Found 2 solutions by josgarithmetic, greenestamps:
Answer by josgarithmetic(39792)
Answer by greenestamps(13327)



(1) The link provided by the other tutor does not seem to me like a particularly good one.

(2) There are hundreds of problems like this on this forum, as well as lessons on how to solve mixture problems algebraically.

(3) And here is a very quick and easy non-algebraic way to solve any two-part mixture problem like this....

Picture the three percentages on a number line: 10, 34, and 50.
In going from 10 towards 50 and stopping at 34, what fraction of the distance do you go? 10 to 50 is a difference of 40; 10 to 34 is a difference of 24; 34 is 24/40 = 3/5 of the way from 10 to 50.
You went 3/5 of the way from 10 to 50; that means 3/5 of the mixture is the 50% solution.

ANSWER: 3/5 of 100ml, or 60ml, of the 50% solution, and the other 40ml of the 10% solution.

CHECK:
60(.50)+40(.10( = 30+4 = 34
100(.34) = 34